How many times a wheel of radius 28 cm must rotate to go 352 m? (Take π = 22/7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Radius of the wheel = 28 cm
Circumference of the wheel = 2πr
= 2 × 22/7 × 28
= 2 × 22 × 4 = 176 cm
Now we have to find the number of rotation of the wheel,
= Total distance to be covered/ circumference of wheel = 352 m/176 cm
= 35200 cm/ 176 cm
= 200

Thanks